# SOFIA/HAWC+ FAR-INFRARED POLARIMETRIC LARGE-AREA CMZ EXPLORATION (FIREPLACE) SURVEY III: FULL SURVEY DATA SET

### v2

@@ 5 THE MAGNETIC FIELDS OF PROMINENT CMZ MOLECULAR CLOUDS
We have argued the FIREPLACE results predominantly reveal the magnetic field local to the CMZ, especially in bright emission regions. In this section, we compare the FIREPLACE 214 µm results with previous polarimetric observations of individual molecular clouds. Furthermore, the increased sensitivity of the FIREPLACE survey enables a determination of how the magnetic field local to the CMZ varies in the peripheries of these clouds than previous studies. Here we present an overview of the magnetic field configurations obtained for a set of prominent CMZ molecular clouds, as indicated in Figure 15.
In the following discussion Figures 16 - 20 are used to motivate the discussion of the magnetic fields local to the distinct clouds. These figures all show how the FIREPLACE magnetic field pseudovectors compare to previous polarimetric observations of these clouds. In general, our new 214 µm observations find a magnetic field that is consistent with these previous studies. Cases where this is not the case are discussed in the relevant subsections.
@@
M0.253+0.016, known colloquially as “The Brick,” lies to the Galactic West of the Dust Ridge and is near the arched filaments that are proximal to the Radio Arc NTFs (Lang et al., 1999; Yusef-Zadeh & Morris, 1987b). The location of the Brick in the GC can be observed in Figure 15. A zoom-in view of the Brick’s total intensity emission obtained from Herschel at 250 µm is shown in Figure 16. Overlaid on this total intensity distribution are the magnetic field pseudovectors derived from our FIREPLACE observations (blue line segments), the magnetic field pseudovectors derived from 350 µm CSO observations (red line segments, Dotson et al., 2010; Pillai et al., 2015), and the magnetic field pseudovectors derived from 850 µm JCMT observations (cyan line segments, Lu et al., 2023). We obtain 50 magnetic field pseudovectors corresponding to the Brick, extending into fainter emission regions than the CSO or JCMT observations. The FIREPLACE pseudovectors are observed to trace the total intensity morphology of the Brick, generally following the extended curvature of the cloud. Our FIREPLACE magnetic field orientations also agree with those derived from the HAWC+ CNM mode observations of this cloud as seen in Figure 6.
@@
The ordered magnetic field obtained from the CSO observations led Pillai et al. (2015) to postulate that the magnetic pressure exceeds the turbulent pressure in this cloud. If true, this indicates that the magnetic field is more dynamically important than turbulence and plays a more significant role in shaping the evolution of the Brick. The similarly ordered magnetic field derived from our FIREPLACE observations support this possibility, and we furthermore observe that the ordered field extends into fainter regions of the Brick than was measured by Pillai et al. (2015). Additionally, HCO+ absorption filaments observed within the Brick also appear to trace the magnetic field orientation within the cloud (Bally et al., 2014), which indicates that the density structures within the cloud could be shaped by the internal magnetic field geometry.
@@
Pillai et al. (2015) argue that the curvature of the magnetic field seen towards the Brick implies the presence of a shock. A recent study of the kinematics and structure of the Brick reveals that the cloud is likely a conglomeration of distinct velocity components along the line-of-sight (Henshaw et al., 2019, 2022), which could result from the Brick consisting of distinct molecular structures that, although superimposed along the line-of-sight, are at different locations in the gravitational potential of the CMZ.
@@
There is also evidence of feedback in the Brick given the presence of an expanding bubble likely driven by the stellar wind of a high mass star (Henshaw et al., 2022). The extent of this wind-blown bubble is marked with a red dashed circle in Figure 16. Feedback shocks resulting from the expansion of this wind-blown bubble could generate the shocked gas and explain the ordered and curved magnetic field of the Brick. In fact, inspection of the magnetic field pseudovectors that coincide with the periphery of the bubble reveals that they largely trace the periphery of the bubble, particularly in high emission regions coincident with the core of the Brick molecular cloud.
@@
The polarization fractions from the Brick are observed to increase towards the periphery of the cloud (5%), whereas the fractional polarizations in the core of the cloud are generally lower (1%). Increased polarization fraction towards the periphery of the Brick is consistent with the CMZ-wide trend of increased depolarization with higher emission as shown in Figure 13.
@@
M0.145-0.086 (Straw), M0.106-0.082 (Sticks), and M0.068-0.075 (Stone) together comprise the cloud complex known as the “Three Little Pigs” (TLP, Battersby et al., 2020). These clouds are labelled and shown in Figure 17. This complex is oriented nearly parallel to the Galactic plane. The clouds all have similar velocities which indicates they are associated with each other and not simply close in projection. The TLP is located 10′ in projection from Sgr A as can be seen in Figure 15.
@@
Previous polarimetric observations of the TLP clouds yielded preliminary evidence of a bending magnetic field within the Stone cloud in the TLP and on the periphery of this cloud closest to Sgr A (Chuss et al., 2003a, b; Lu et al., 2023). the Straw and the Sticks clouds, conversely, were observed to have magnetic fields that are oriented largely perpendicular to the Galactic plane and do not exhibit as much warping as seen in the Stone cloud. This result is also observed in the arcminute-resolution ACTPol results presented in Guan et al. (2021). These previous polarimetric observations were either only sensitive to the brightest regions of the clouds in the TLP or were limited by lower angular resolutions, and so it was unclear from those observations how the field behaves between the clouds of the TLP and the outer periphery of the TLP complex.
Our new FIREPLACE observations provide over 100 magnetic field pseudovectors corresponding to the TLP that uncover the magnetic field between the TLP clouds and in the lower emission regions at the periphery of the cloud complex. The FIREPLACE 214 µm dust polarimetry observations recover magnetic field orientations in the cores of the Straw, the Sticks, and the Stone clouds similar to those derived from the CSO and JCMT observations.
@@
The magnetic field surrounding the Straw cloud derived from FIREPLACE is perpendicular to the Galactic plane at lower Galactic latitudes (-0.1 and below in the figure) whereas the magnetic field above the cloud is angled almost parallel to the Galactic plane. The field above and below the Sticks cloud is oriented perpendicular to the Galactic plane and only becomes warped for lines of sight where the emission exceeds MJy sr. The field surrounding the Stone cloud exhibits the same curve as is observed in the core of the cloud from the previous CSO and JCMT observations.
@@
The systematic differences in magnetic field orientation seen above and below the Straw cloud coincide with the low-longitude outer extent of the M0.20-0.030 shell likely originating from the Quintuplet cluster (Butterfield et al., 2022, 2018). Butterfield et al. (2018) argued that the TLP clouds could be interacting with the M0.20-0.033 shell. The magnetic field pseudovectors from FIREPLACE do appear to be impacted by the presence of the shell: changing from a systematically perpendicular (with respect to the Galactic plane) magnetic field orientation outside the shell to parallel or rotated within the shell.
@@
The impact of the shell on the magnetic field pseudovectors is also observed in higher emission regions of the Straw cloud where the core of the cloud coincides with the outer extent of the shell. This result corroborates the argument of Butterfield et al. (2018) of an interaction between the shell and the TLP complex; indicating that at least the Straw cloud is likely impacted by the presence of the shell. A similar argument is made in Lu et al. (2023) based on their JCMT observations of this cloud.
@@
Neither the Sticks nor the Stone clouds show any evidence of interaction with the M0.20-0.030 shell. However, the field coinciding with and surrounding the Stone cloud is noticeably more warped than that of the Sticks cloud. The Stone cloud in particular is too far in projection from the M0.20-0.033 shell for the curvature of the magnetic field in that region to be explained by interaction with that shell. Instead, the curvature could be the deformation of an initially vertical magnetic field induced by the orbital motion of the TLP cloud complex as it moves through the field. Additional analysis is needed to determine the connection between the properties of the TLP clouds and the larger field structure.
@@
An explanation for the differences in the magnetic field geometries observed between the Straw, the Sticks, and the Stone cloud could be differences in mass between the different clouds. There is a distinct brightness gradient in the emission from the clouds in the TLP (where the brightness here is proportional to the column density), with the Straw cloud being the faintest and M0.068-0.75 being the brightest. The differences in mass for the clouds in the TLP could indicate changes in the magnetic critical parameter which quantifies the relative importance of the magnetic field and self-gravity within these clouds (Guerra et al., 2023).
@@
A difference in the critical parameter between the clouds could indicate changes in the tendency of the clouds to contract and form stars (Nakano & Nakamura, 1978). This possibility is consistent with recent star formation rate (SFR) estimates obtained by the CMZoom survey that found an SFR of 6.43.8 M yr for the Stone cloud compared to 0.40.3 M yr for the Straw cloud (Hatchfield et al., 2023). The higher SFR of the Stone cloud increases the likelihood of there being outflowing winds within this cloud, similar to the bubble observed by Henshaw et al. (2019) within the Brick. Follow-up work is necessary to determine whether it is the orbital motion of the TLP or differences in the presence of stellar winds between the clouds that best explains the differences in the magnetic field seen for the Sticks and the Stone cloud.
@@
As with the Brick we observe increased polarization fractions towards the periphery of the TLP cloud complex. The highest polarization fractions observed for this cloud are 5%, with regions in the center of the cloud having generally lower polarization fractions (1%) than what is observed in the fainter peripheries of the cloud complex.
5.3 Clouds Surrounding Sgr A
@@
There are multiple clouds that are within 5′ in projection from Sgr A. The prominent clouds that are within this region and shown in Figure 15 are M0.001-0.058 (50 km s cloud), M359.948-0.052 (Circum-nuclear Disk), CO 0.02-0.02 (as labeled in Lu et al. 2023), and M359.889-0.093 (20 km s cloud). We analyze these clouds as a group and they are presented in Figures 18 and 19.
5.3.1 M0.001-0.058 (50 km s Cloud)
@@
The 50 km s cloud is located to the Galactic Southeast of Sgr A and is contained within the greater Sgr A complex. The location of the 50 km s cloud is shown in Figure 15.
@@
We obtain 50 magnetic field pseudovectors from our 214 µm observations and are largely oriented perpendicular to the Galactic plane. There are deviations from this perpendicular field orientation. For example, there is a dense dust ridge that lies along the Eastern periphery of the Sgr A East supernova remnant (SNR Serabyn et al., 1992) at = 359.967, = -0.067. The magnetic field coincident with this dust ridge traces the orientation of this ridge, which is rotated by 45 relative to the Galactic plane. This dust ridge in the 50 km s cloud is thought to be compressed by a SNR (Ehlerová et al., 2022; Novak et al., 2000; Serabyn et al., 1992), which can explain why the magnetic field is oriented along the ridge.
@@
The magnetic field observed in the 50 km s cloud changes direction at Southern Galactic latitudes where the intensity decreases and the previous CSO and JCMT observations lack sensitivity. This is particularly noticeable at = 359.97, = -0.117 where the field subtends an angle of 60 with respect to the Galactic plane. The fractional polarizations are also quite large in this region (generally 5%). The magnetic field orientation also systematically changes from being generally oriented perpendicular to the Galactic plane to the East of the 50 km s cloud to being parallel to the Galactic plane to the West of the 50 km s cloud.
@@
The Circum-nuclear Disk (CND) is a set of clouds located within 5 pc of Sgr A. The CND has an inclination of 70 relative to the plane of the sky and an inner radius of pc (Guerra et al., 2023; Lau et al., 2013). The CND is labeled in Figure 18.
@@
Continuum Stokes imaging with HAWC+ at 19.7, 31.5, 37.1, and 53 µm show a very clear set of clouds that are gravitationally related (either orbiting or infalling) to Sgr A* (Lau et al., 2013; Guerra et al., 2023) and are experiencing gravitational shear. At far-infrared wavelengths Molinari et al. (2011), the structure of the CND is only faintly apparent, with the emission being dominated by more extended, cooler CMZ dust, which is located external to the central 5 pc. The 214 µm Stokes maps are consistent with these images.
In polarimetry, the CND has been measured at 53 µm at 5″ resolution. These pseudovectors show a field that is highly aligned with the dust streamer morphology (Guerra et al., 2023) who argue for a steady-state shear flow leading to a magnetic field that traces the morphology of the CND arms. They estimate a milligauss magnetic field strength for the fields in the CND using a modified version of the Davis-Chandrasekhar-Fermi technique (Davis, 1951; Chandrasekhar & Fermi, 1953) that takes shear into account.
The FIREPLACE dataset contains 50 Nyquist-sampled magnetic field pseudovectors in the vicinity of the CND. In contrast to the shorter wavelength observations, the inferred magnetic field is quite ordered and is largely oriented parallel to the projected long axis of the CND, which itself is oriented at an angle of about 20 degrees with respect to the Galactic plane.
@@
The FIREPLACE 214 µm data is consistent with the geometry measured at 350 µm at 20″ resolution (Novak et al., 2000). Parts of the CND were included in the SCUBA-2/Pol-2 map at 850 µm and 20″ resolution (Lu et al., 2023). There is general agreement with these vectors in the Northeastern part of this region; however, to the south, the 850 µm vectors do not agree with the shorter wavelengths. This could be an intrinsic variation in the source polarization at these different wavelengths or it could be a result of differences in spatial filtering from the different observing modes (OTFMAP map for this study compared to CNM mode for the CSO and Daisy mode for the JCMT (Lu et al., 2023)). However, our FIREPLACE results agree with the HAWC+ CNM mode observations of this cloud as seen in Figure 6. Therefore the difference is likely a result of changes in the polarization over the different wavelengths. Previous 850 µm polarization measurements have been made of the CND region (Hsieh et al., 2018), in which the authors claimed to detect a sheared geometry of the field.
@@
The differences in morphology and polarization orientation from 53 to 214 µm (and longer wavelengths) are most likely due to two cloud components superposed along the line-of-sight the warm dust local to the streamers of the CND and the cooler dust component in the wider CMZ. The former contains magnetic fields that dynamically connected to the sheared flow of the inner 5pc; the latter are likely shaped by the dynamics of the greater CMZ.
@@
Cloud CO 0.02-0.02 is only 4′ away from the CND in projection to the Northeast of the CND as can be seen in Figure 18. The FIREPLACE observations reveal 20 magnetic field pseudovectors coinciding with the CO 0.02-0.02 cloud that are generally oriented perpendicular to the Galactic plane; however, this orientation does not agree with what is observed in previous CSO and JCMT observations of this cloud. These pseudovectors differ by a 30 angle with respect to the FIREPLACE pseudovectors. In addition, there are instances where the CSO and JCMT observations disagree, which could indicate complexity in this region of the cloud.
@@
The disagreement in magnetic field orientation could be a result of different spatial filtering between our FIREPLACE observations and that of the CSO and JCMT observations. In addition, we note that the CO 0.02-0.02 cloud falls within the region of elevated and variance observed in Figure 5 from our FIREPLACE observations. The angle differences observed for this cloud could therefore indicate remaining systematic uncertainties in our FIREPLACE observations. This explanation is further justified by the fact that the CSO and JCMT observations are in general agreement over most of this cloud.
5.3.4 M359.889-0.093 (20 km s Cloud)
@@
M359.889-0.093, known colloquially as the 20 km s cloud, is located 10′ in projection from Sgr A to the Southwest as shown in Figure 15. A zoom-in view of the 20 km s cloud is shown in Figure 19. The orange arrow in Figure 19 indicates the direction to Sgr A.
@@
Our SOFIA/HAWC+ observations yield 100 distinct pseudovectors that span the full extent of the 20 km s cloud. The field within the cloud is predominantly oriented parallel to the Galactic plane. The field within the cloud is seemingly preserved above and below the cloud in Galactic latitude, being also largely oriented parallel to the Galactic plane.

### v3

@@ 5 THE MAGNETIC FIELDS OF PROMINENT CMZ MOLECULAR CLOUDS
We have argued the FIREPLACE results predominantly reveal the magnetic field local to the CMZ, especially in bright emission regions (Section 4.2). In this section, we compare the FIREPLACE 214 µm results with previous polarimetric observations of individual molecular clouds. Furthermore, the increased sensitivity of the FIREPLACE survey enables a better determination of how the magnetic field local to the CMZ varies in the peripheries of these clouds than previous studies. Here we present an overview of the magnetic field configurations obtained for a set of prominent CMZ molecular clouds, as indicated in Figure 16.
In the following discussion, Figures 17 - 21 are used to motivate the discussion of the magnetic fields local to the distinct clouds. These figures show how the FIREPLACE magnetic field pseudovectors compare to previous polarimetric observations of these clouds. In general, our new 214 µm observations find a magnetic field that is consistent with these previous studies. Instances where this is not the case are discussed in the relevant subsections.
@@
M0.253+0.016, known colloquially as the “Brick,” lies to the Galactic Northeast of Sgr A∗ and is near the arched filaments that are proximal to the Radio Arc NTFs (Yusef-Zadeh & Morris, 1987b; Lang et al., 1999). The location of the Brick in the GC can be observed in Figure 16. A zoom-in view of the Brick’s total intensity emission obtained from Herschel at 250 µm is shown in Figure 17. Overlaid on this total intensity distribution are the magnetic field pseudovectors derived from our FIREPLACE observations (blue line segments), the magnetic field pseudovectors derived from 350 µm CSO observations (red line segments, Dotson et al., 2010; Pillai et al., 2015), and the magnetic field pseudovectors derived from 850 µm JCMT observations (cyan line segments, Lu et al., 2024). We obtain 50 magnetic field pseudovectors corresponding to the Brick, extending into fainter emission regions than the previous CSO or JCMT observations. The FIREPLACE pseudovectors are observed to trace the total intensity morphology of the Brick, generally following the extended curvature of the cloud. Our FIREPLACE magnetic field orientations also agree with those derived from previous HAWC+ CNM mode observations of this cloud, as seen in Figure 6.
@@
The ordered magnetic field obtained from the CSO observations led Pillai et al. (2015) to postulate that the magnetic pressure exceeds the turbulent pressure in this cloud. If true, this indicates that the magnetic field is more dynamically important than turbulence and plays a more significant role in shaping the morphology of the Brick. The similarly ordered magnetic field derived from our FIREPLACE observations support this possibility, and we furthermore observe that the ordered field extends into fainter regions of the Brick than was measured by Pillai et al. (2015). Additionally, HCO+ absorption filaments observed within the Brick also appear to trace the magnetic field orientation within the cloud (Bally et al., 2014), which indicates that the density structures within the cloud could be shaped by the internal magnetic field geometry.
@@
Pillai et al. (2015) argue that the curvature of the magnetic field seen towards the Brick implies the presence of a shock. A recent study of the kinematics and structure of the Brick implies that the cloud is likely a conglomeration of distinct velocity components along the line-of-sight (Henshaw et al., 2019, 2022). The Brick could therefore be multiple distinct molecular structures that, although superimposed along the line-of-sight, are at different locations in the gravitational potential of the CMZ.
@@
There is also evidence of feedback in the Brick given the presence of an expanding bubble likely driven by the stellar wind of a high mass star (Henshaw et al., 2022). The expanding bubble could be driving the shock observed in Pillai et al. (2015). The extent of this wind-blown bubble (a radius of 1.3 pc, Henshaw et al., 2022) is marked with a red dashed circle in Figure 17. Feedback shocks resulting from the expansion of this wind-blown bubble could generate the shocked gas and explain the ordered and curved magnetic field of the Brick. In fact, inspection of the FIREPLACE magnetic field pseudovectors that coincide with the periphery of the bubble reveals that they largely trace the periphery of the bubble, particularly in high emission regions coincident with the core of the Brick molecular cloud.
@@
The fractional polarizations North of the Brick are generally 5%, whereas the fractional polarizations coinciding with the Brick range from 5% to 1%. The fractional polarization within the Brick is observed to steadily decrease from North to South in Galactic Latitude. Outside of the Brick we observe fractional polarizations that are generally 5% (like for the pseudovectors observed to the Southwest of the Brick). The large fractional polarization observed external to the Brick, coupled with the generally lower fractional polarizations internal to the cloud, is consistent with the CMZ-wide trend of increased depolarization with higher emission as shown in Figure 13.
@@
The M0.145-0.086 (Straw), M0.106-0.082 (Sticks), and M0.068-0.075 (Stone) comprise the cloud complex known as the “Three Little Pigs” (TLP, Battersby et al., 2020). These clouds are labeled and shown in Figure 18. The extent of this complex is oriented nearly parallel to the Galactic plane. The clouds all have similar velocities (50 km s-1 e.g., Kruijssen et al., 2015; Butterfield et al., 2018). This indicates they are likely dynamically associated with each other and not simply close in projection. The TLP is located 10′ in projection from Sgr A∗ as can be seen in Figure 16.
@@
Previous polarimetric observations of the TLP clouds yielded preliminary evidence of a curved magnetic field within the Stone cloud in the TLP. This curvature occurs on the periphery of this cloud closest to Sgr A∗ (Chuss et al., 2003a, b; Lu et al., 2024). The Straw and the Sticks clouds, conversely, were observed to have magnetic fields that are oriented largely perpendicular to the Galactic plane and do not exhibit as much warping as seen in the Stone cloud. The perpendicular orientation of the magnetic field near the Straw cloud is particularly interesting since this cloud coincides with the Radio Arc NTF (e.g. Yusef-Zadeh et al., 1984; Yusef-Zadeh & Morris, 1987c; Paré et al., 2019), as can be seen in Figure 16. This perpendicular field is parallel to the orientation of the filaments in the Radio Arc, which could indicate a connection between the field inferred from dust polarization and the field traced by the synchrotron emission of the Radio Arc filaments.
The difference in magnetic field curvature between the TLP clouds was previously observed in the arcminute-resolution ACTPol results presented in Guan et al. (2021). Our results are consistent with these previous observations but show more detail thanks to our 19.6″ resolution. Our new FIREPLACE observations provide over 100 magnetic field pseudovectors in the vicinity of the TLP, including most of the extended, lower intensity regions. In the cores of the Straw, the Sticks, and the Stone clouds the FIREPLACE pseudovectors have similar orientations to those derived from the CSO and JCMT observations.
@@
The FIREPLACE magnetic field to the South of the Straw cloud is perpendicular to the Galactic plane at lower Galactic latitudes (-0.1∘ and below in the figure). Conversely, the magnetic field above the cloud is oriented almost parallel to the Galactic plane. The field above and below the Sticks cloud is oriented perpendicular to the Galactic plane and only becomes warped for lines-of-sight where the emission exceeds MJy sr-1. The field surrounding the Stone cloud exhibits the same curvature as is observed in the core of the cloud from the previous CSO and JCMT observations.
@@
The systematic differences in the magnetic field orientation seen above and below the Straw cloud coincide with the low-latitude outer extent of the M0.20-0.033 shell likely originating from the Quintuplet cluster (Butterfield et al., 2018, 2022) as marked in Figure 18. They argue that the TLP clouds could be interacting with the M0.20-0.033 shell. The magnetic field pseudovectors from FIREPLACE do appear to be impacted by the presence of the shell: changing from a systematically perpendicular (with respect to the Galactic plane) magnetic field orientation outside the shell to parallel or rotated within the shell.
@@
The impact of the shell on the magnetic field pseudovectors is also observed in higher emission regions of the Straw cloud, where the core of the cloud coincides with the outer extent of the shell. This result corroborates the argument of Butterfield et al. (2018, 2022) of an interaction between the shell and the TLP complex; indicating that at least the Straw cloud is likely impacted by the presence of the shell. A similar argument is made in Lu et al. (2024) based on their JCMT observations of this cloud.
@@
Neither the Sticks nor the Stone clouds show any evidence of interaction with the M0.20-0.033 shell in our observations. However, the field coinciding with and surrounding the Stone cloud is noticeably more warped than that of the Sticks cloud. The Stone cloud in particular is too far in projection from the M0.20-0.033 shell for the curvature of the magnetic field in that region to be explained by interaction with that shell. Instead, the curvature could be the deformation of an initially vertical magnetic field induced by the orbital motion of the TLP cloud complex as it moves through the field.
@@
An explanation for the differences in the magnetic field geometries observed between the Straw, Sticks, and Stone clouds could be due to differences in column density, as can be seen from the Herschel column density maps of the GC (Molinari et al., 2011), with the Straw cloud being the least dense and the Stone being the most dense. This column density gradient could indicate changes in the magnetic critical parameter which quantifies the relative importance of the magnetic field and self-gravity within these clouds (e.g., Nakano & Nakamura, 1978).
@@
A difference in the magnetic critical parameter between the clouds could indicate changes in the tendency of the clouds to contract and form stars (Nakano & Nakamura, 1978). This possibility is consistent with recent star formation rate (SFR) estimates obtained by the CMZoom survey. They found an SFR of 6.43.8 10-3M⊙ yr-1 for the Stone cloud compared to 0.40.3 10-3M⊙ yr-1 for the Straw cloud (Hatchfield et al., 2024). The higher SFR of the Stone cloud increases the likelihood of there being outflowing winds within this cloud. However, it remains unclear what the cause for the different magnetic field distribution could be. Follow-up work is necessary to determine whether it is the orbital motion of the TLP or differences in the presence of stellar winds between the clouds that best explains the differences in the magnetic field seen for the Sticks and the Stone cloud.
@@
As with the Brick, we observe increased polarization fractions towards the periphery of the TLP cloud complex. The highest polarization fractions observed for this cloud are 5%, with regions in the center of the cloud having generally lower polarization fractions (1%) than what is observed in the fainter peripheries of the cloud complex.
5.3 Clouds Surrounding Sgr A∗
@@
There are multiple clouds that are within 5′ in projection from Sgr A∗ (Figure 16). These prominent clouds are M0.001-0.058 (50 km s-1 cloud), M359.948-0.052 (Circum-nuclear Disk), CO 0.02-0.02 (as labeled in Lu et al. 2024), and M359.889-0.093 (20 km s-1 cloud). Zoom-in views of these clouds are presented in Figures 19 and 20.
5.3.1 M0.001-0.058 (50 km s-1 Cloud)
@@
The 50 km s-1 cloud is located to the Galactic Southeast of Sgr A∗ and is contained within the greater Sgr A complex (Figure 16).
@@
We obtain 50 magnetic field pseudovectors from our 214 µm observations coinciding with the 50 km s-1 cloud as seen in Figure 19. These pseudovectors are largely oriented perpendicular to the Galactic plane. However, there are deviations from this perpendicular field orientation. For example, there is a dense dust ridge within the 50 km s-1 cloud at = 359.967∘, = -0.067∘. The magnetic field coincident with this dust ridge traces the orientation of this ridge, which is rotated by 45∘ relative to the Galactic plane. This dust ridge in the 50 km s-1 cloud is thought to be compressed by the nearby Sgr A East supernova remnant, the extent of which is marked by the dashed yellow circle in Figure 19 (Serabyn et al., 1992; Novak et al., 2000; Ehlerová et al., 2022), which can explain why the magnetic field is oriented along the ridge. The dust ridge traces the edge of the supernova remnant, meaning the FIREPLACE magnetic field is largely azimuthal with respect to the supernova remnant. A similar, largely azimuthal, magnetic field morphology is observed in the periphery of the Ring structure studied in FIREPLACE II.
@@
The magnetic field observed in the 50 km s-1 cloud changes direction at Southern Galactic latitudes where the intensity decreases and the previous CSO and JCMT observations lack sensitivity. This is particularly noticeable at = 359.97∘, = -0.117∘ where the field subtends an angle of 60∘ with respect to the Galactic plane. The fractional polarizations are also quite large in this region (generally 5%). The magnetic field orientation also systematically changes from being generally oriented perpendicular to the Galactic plane to the East of the 50 km s-1 cloud to being parallel to the Galactic plane to the West of the 50 km s-1 cloud.
@@
The Circum-nuclear Disk (CND) is located within 5 pc of the super massive black hole Sgr A∗ (labeled in Figure 19). The CND has an inclination of 70∘ relative to the plane of the sky and an inner radius of pc (Lau et al., 2013; Guerra et al., 2023).
@@
Images by SOFIA/FORCAST at 19.7, 31.5, and 37.1, and SOFIA/HAWC+ images at 53 µm show a clear indication that the CND is experiencing gravitational shear (Lau et al., 2013; Guerra et al., 2023). At far-infrared wavelengths (250 µm, Molinari et al., 2011), the structure of the CND is only faintly apparent, with the emission being dominated by more extended, cooler CMZ dust, which is located external to the central 5 pc. The FIREPLACE 214 µm Stokes maps are consistent with these images.
@@
The FIREPLACE dataset contains 50 Nyquist-sampled magnetic field pseudovectors in the vicinity of the CND. The FIREPLACE 214 µm data is consistent with the geometry measured at 350 µm at 20″ resolution (red pseudovectors in Figure 19, Novak et al., 2000). Parts of the CND were included in the SCUBA-2/Pol-2 map at 850 µm and 20″ resolution (cyan pseudovectors in Figure 19; Lu et al., 2024). There is general agreement with these vectors in the Northeastern part of the CND region; however, to the South, the 850 µm vectors do not agree with the shorter wavelengths. This disagreement could be an intrinsic variation in the source polarization at these different wavelengths. Alternatively, it could be a result of differences in spatial filtering from the different observing modes (OTFMAP mode for this study compared to CNM mode for the CSO and Daisy mode for the JCMT (Lu et al., 2024)). However, our FIREPLACE results agree with the HAWC+ CNM mode observations of this cloud as seen in Figure 6. Therefore the difference is likely a result of changes in the polarization over the different wavelengths. Previous 850 µm polarization measurements have been made of the CND region (Hsieh et al., 2018), in which the authors claimed to detect a sheared geometry of the field.
The claim of a sheared field was supported by more recent observations of the CND (Guerra et al., 2023). The pseudovectors from these observations show a field that is highly aligned with the dust streamer morphology in Guerra et al. (2023), who argue for a steady-state shear flow leading to a magnetic field that traces the morphology of the CND arms (alternatively referred to as “wings” in other CND studies e.g., Zhao et al., 2016). They estimate a mG magnetic field strength in the CND using a modified version of the Davis-Chandrasekhar-Fermi technique that accounts for shear (Davis, 1951; Chandrasekhar & Fermi, 1953).
@@
The differences in morphology and polarization orientation from 53 to 214 µm (and longer wavelengths) are most likely due to two cloud components superposed along the line-of-sight: the warm dust local to the streamers of the CND and the cooler dust component in the wider CMZ. The former contains magnetic fields that are dynamically connected to the sheared flow of the inner 5 pc; the latter are likely shaped by the dynamics of the greater CMZ.
@@
Cloud CO 0.02-0.02 is only 4′ away from the CND in projection to the Northeast of the CND as can be seen in Figure 19. The FIREPLACE observations reveal 20 magnetic field pseudovectors coinciding with the CO 0.02-0.02 cloud. These pseudovectors are generally oriented perpendicular to the Galactic plane; however, this orientation does not agree with what is observed in previous CSO and JCMT observations of this cloud (Lu et al., 2024). These pseudovectors differ by a 30∘ angle with respect to the FIREPLACE pseudovectors. In addition, there are instances where the CSO and JCMT observations disagree, which could indicate magnetic field complexity in this region of the cloud.
@@
The disagreement in magnetic field orientation could be a result of different spatial filtering between our FIREPLACE observations and that of the CSO and JCMT observations. In addition, we note that the CO 0.02-0.02 cloud falls within the region of elevated and variance observed in Figure 5 from our FIREPLACE observations. The angle differences observed for this cloud could therefore indicate remaining systematic uncertainties in our FIREPLACE observations. This explanation is further supported by the fact that the CSO and JCMT observations are in general agreement over most of this cloud.
5.3.4 M359.889-0.093 (20 km s-1 Cloud)
@@
M359.889-0.093, known colloquially as the 20 km s-1 cloud, is located 10′ in projection from Sgr A∗ to the Southwest as shown in Figure 16. A zoom-in view of the 20 km s-1 cloud is shown in Figure 20. The orange arrow in Figure 20 indicates the direction to Sgr A∗.
@@
Our SOFIA/HAWC+ observations yield 100 distinct pseudovectors that span the full extent of the 20 km s-1 cloud. The field within the cloud is predominantly oriented parallel to the Galactic plane. The field above and below the cloud in Galactic latitude is also largely oriented parallel to the Galactic plane.